# GL1-02-06-24-2-supplement

### v2

@@ 10 Conclusion
This paper develops a generalized Difference-in-Difference model for discrete, possibly underreported outcomes in repeated cross-sectional or panel data. The models are based on the continuous CiC model of Athey and Imbens (2006). For the case without misreporting and covariates, we propose a parametric model that generalizes standard Probit or Logit DiD models which only accommodate mean shifts. When covariates are present (but outcomes are reported correctly), we relax the distributional assumption on the unobservables, and provide a point identification result for a semiparametric index model. The result is complementary to Athey and Imbens (2006) and offers a direct pathway to estimation of the structural model and corresponding treatment effects.
@@
In the second step, we extend the CiC model for discrete outcomes to partially observable outcomes. Special features of this extended model are that it allows the underreporting behavior to be correlated with the outcome of interest and that underreporting may depend on the treatment status itself. We provide point identification conditions for a semiparametric version of the model based on a real analytic extrapolation assumption, and suggest a flexible estimator for a parametric version of the model.

### v3

@@ 6 Conclusion
This paper develops a Difference-in-Differences model for discrete, possibly underreported ordered outcomes in repeated cross-sectional or panel data. The models are based on the continuous CiC model of Athey and Imbens (2006). We extend this CiC model for discrete outcomes to partially observable outcomes. Special features of this extended model are that it allows the underreporting behavior to be correlated with the outcome of interest and that underreporting may depend on the treatment status itself. We provide point identification conditions for a semiparametric version of the model based on a real analytic extrapolation assumption, and suggest a flexible estimator for a parametric version of the model.